OpenStudy (anonymous):

find the values of x at which the tangent line to the graph is parallel to the line y=x. y= (x^2+1)/(x+1)

OpenStudy (anonymous):

wait would the derivative be 1? meaning it would be a line meaning no tan line?

OpenStudy (anonymous):

y-x has a slope of 1 so if you differentiate y= (x^2+1)/(x+1) and equate it to and then solve for x you will get the required values of x

OpenStudy (anonymous):

* equate it to 1

OpenStudy (anonymous):

well for the derivative I got the answer: x^2+2x+1/(x+1)^2 which would be 1 because it is the same top and bottom?

OpenStudy (anonymous):

the answer in the back says none, i'm just tryind to figure out why.

OpenStudy (anonymous):

i'll just check it out again

OpenStudy (anonymous):

ok so if you set the derivative to 1 then both sides would be equal to each other. leaving nothing. is that how that would be done?

OpenStudy (anonymous):

whats the matter with me today - sorry - the derivative y' = 1 is a line of slope 1 which is a line parallel to y=x. two parallel liones never intersect so there are no solutions

OpenStudy (anonymous):

how wud you do a problem like this when it asks for it to be perpendicular? its y=(x+3)/(x+2)

OpenStudy (anonymous):

perpendicular to y=x?

OpenStudy (anonymous):

yes. i got the derivative as -1/(x+2)^2

OpenStudy (anonymous):

yes derivative is correct

OpenStudy (anonymous):

so what would I be setting it equal to for a perpendicular line? would be be the opposite reciprical?

OpenStudy (anonymous):

the derivative gives the slope of the tangent to the curve

OpenStudy (anonymous):

so y = x always has the slope of 1? so would you set it to -1?

OpenStudy (anonymous):

Sorry, I'm so confused.

OpenStudy (anonymous):

the derivative -1/(x+2)^2 gives us the slope of all points x on the curve (x+3)/(x+ 2) eg at x = 1 the slope is -1 /(1+2)^2 = -1/8. we are required to find values of x which will give a line perpendicular to y = x , right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok the slope of a line perpendicular to a line of slope m = -1/m. in case of y=s the slope is -1 so -1/(x+2)^2 = -1 -(x+2)^2 = -1 (x+2)^2 = 1 x^2 + 4x + 3 = 0 solve this to get your answers

OpenStudy (anonymous):

x = -1 , -3

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

yes that is what I got. Thank you for helping me